OpenStudy (anonymous):

please help with this problem 3x^3+4x^2=27x+36

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you have a cubic equation to solve?

OpenStudy (anonymous):

yes can you please explain all the steps for me

OpenStudy (misty1212):

there is a formula for solving cubes, but it is almost incomprehensible start with \[3x^3+4x^2-27x-36=0\] and then make an educated guess , which is really all you ca ndo

OpenStudy (misty1212):

i guess 3, since 1 and -1 do not work shall we try it?

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

i;ll be damned, it works!

OpenStudy (misty1212):

got it in the first try so this is \[(x-3)(\text{ some quadratic})\]which you can find using synthetic division or long division (ugh) or thinking (or cheating)

OpenStudy (misty1212):

i can't write synthetic division here, but i can cheat or think you pick

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

tnx

OpenStudy (johnn.):

Let's solve your equation step-by-step. 3x3+4x2=27x+36 Step 1: Subtract 27x+36 from both sides. 3x3+4x2−(27x+36)=27x+36−(27x+36) 3x3+4x2−27x−36=0 Step 2: Factor left side of equation. (x+3)(x−3)(3x+4)=0 Step 3: Set factors equal to 0. x+3=0 or x−3=0 or 3x+4=0 x=−3 or x=3 or x=−43 Answer: x=−3 or x=3 or x=−43